su ahion A bdl is dropped From o heightnd cnuysh rebounds to 2the distance tals a) Whod ithe htad aistance vesed by the ball when comes to rest? b) Whotthe tbital disa splacene
Answer #1

a)

The height after first rebound = 120*2/3 = 80

second rebound = 80*2/3 = 160/3 and so on

The heights are as follows....120,120(2/3), 120(2/3)^2, 120(2/3)^3........

It is an infinte geometric progression whose sum is given by Sn = a/(1-r) = 120/(1-2/3) = 120/(1/3) = 120*3 = 360 ft.

b)

Final position of the ball be when it comes to rest

Hence displacement = 120 - 0 = 120 ft